OpenStudy (anonymous):

A heron can travel an average of 400 km in 10h. What is the rate of change of distance ?

OpenStudy (anonymous):

You're looking for speed, which is the rate of change of distance (with respect to time). v = d/t You are given that: d = 400 km = 400000 m t = 10 h = 36000 s Therefore: v = 400 km / 10 h = 40 km/h OR v = 400000 m / 36000 s = 11.1 m/s

OpenStudy (anonymous):

Thanks for the help ! :)

OpenStudy (anonymous):

No problem! :)